# Integral bases, perfect matchings, and the Petersen graph

### v2

@@ Abstract
Let $G=(V,E)$ be a matching-covered graph, denote by $P$ its perfect matching polytope, and by $L$ the integer lattice generated by the integral points in $P$. In this paper, we give short, polyhedral proofs for two difficult results established by Lovász (1987), and by Carvalho, Lucchesi, and Murty (2002) in a series of three papers totaling over 120 pages. More specifically, we prove that $L$ has a lattice basis consisting solely of incidence vectors of some perfect matchings of $G$, $2x\in L$ for all $x\in\operatorname{lin}(P)\cap\mathbb{Z}^{E}$, and if $G$ has no Petersen brick then $L=\operatorname{lin}(P)\cap\mathbb{Z}^{E}$. Our proof avoids major technical aspects of the previous proofs, the most important of these being a characterization of the dual lattice, and a ‘Petersen-brick-sensitive’ ear-decomposition result for matching-covered graphs. This is achieved by a novel study of the facial structure of the polytope $P$ and its relationship with the lattice $L$. Along the way, we give a new polyhedral characterization of the Petersen graph.
@@ 1 Introduction
Let $G=(V,E)$ be a matching-covered graph, that is, a graph where every edge appears in a perfect matching. Denote by $P(G)$ the polytope whose vertices are the incidence vectors of the perfect matchings of $G$, and by $L(G)$ the (integer) lattice generated by the integral points in $P(G)$, i.e., $L(G)$ is the set of all integer linear combinations of $P(G)\cap\{0,1\}^{E}$.
@@
A seminal result in combinatorial optimization due to Edmonds and Johnson [9] is that $P(G)$ can be described by non-negativity inequalities: $x_{e}\geq 0,e\in E$, degree equations: $x(\delta(v))=1,v\in V$, and odd cut inequalities amounting to $x(\delta(U))\geq 1$ for all odd-sized subsets $U\subset V$. Here, $x(A)$ denotes $\sum_{a\in A}x_{a}$. Later Seymour [17] gave a graph-theoretic proof of this result using Tutte’s characterization of perfect matchings [18]. In the same paper, through a long and technical argument, Seymour proved that in a bridgeless cubic graph, the all-$2$s vector belongs to $L(G)$, and in fact the all-$1$s vector belongs to the lattice if $G$ has no Petersen minor, thus establishing weaker variants of the Berge-Fulkerson conjecture and Tutte’s $4$-flow conjecture, respectively [11, 19].
@@
Naddef [16] and Edmonds, Pulleyblank, and Lovász [10] computed the dimension of $P(G)$ based on structural properties of $G$. These lines of questions led to the development of a fascinating area known as matching theory [14, 12]. In a seminal paper of the area, Lovász [13] gave a far-reaching extension of Seymour’s result and characterized the vectors in $L(G)$, and proved that $2x\in L$ for all $x\in\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$, and if $G$ has no Petersen brick then $L(G)=\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$. Here, $\operatorname{lin}(S)$ denotes the linear hull of $S$ for $S\subseteq\mathbb{R}^{E}$. The core of his argument was proving a difficult lemma that characterized the dual lattice for a non-Petersen brick. This lengthy proof drew on key tools from matching theory, including the notions of barrier cuts, 2-separation cuts, and ear decompositions.
@@
In a series of three excellent papers [5, 6, 7], totaling over 120 pages, Carvalho, Lucchesi, and to Murty proved that $G$ admits an ear decomposition that, vaguely speaking, is sensitive to the numbers of bricks and Petersen bricks, and as a key application, they proved that $L(G)$ has a lattice basis $B$ consisting solely of incidence vectors of some perfect matchings of $G$, that is, $B$ is a (linear) basis for $\operatorname{lin}(L(G))$, and every vector in $\operatorname{lin}(L(G))\cap\mathbb{Z}^{E}$ is an integer linear combination of $B$. This in turn answered a question raised by Murty ([15], Problem 7.3).
Main Theorem.
In this paper, we give relatively short, polyhedral proofs of the theorems mentioned due to Lovász, and Carvalho, Lucchesi, and Murty, namely the following combined theorem. A Petersen brick is any graph whose simplification is the Petersen graph.
@@
Let $G=(V,E)$ be a matching-covered graph, let $L:=L(G)$ and $\bar{L}:=\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$. Then $L$ has a lattice basis consisting solely of some perfect matchings of $G$. Furthermore, if $G$ has $p$ Petersen bricks in its tight cut decomposition, then
@@
Above, $[p]$ denotes $\{1,2,\ldots,p\}$ for $p\geq 1$, and $\emptyset$ for $p=0$. Our proof mostly circumvents the lengthy graph-theoretic arguments, and in particular has no dependence on matching theoretic notions such as $2$-separation cuts, braces, removable edges, and ear decompositions, and minimal dependence on the notions of barriers and bricks. It also eliminates the need to study the dual lattice through matching-integral vectors. Instead, our proof is based on polyhedral, or otherwise polyhedrally-driven graph-theoretic arguments that study the relationship between the lattice and the facial structure of the polytope.
@@
Let $G=(V,E)$ be a Petersen-free matching-covered graph. Then $\operatorname{lin}(P(G))$ has an integral basis consisting solely of the incidence vectors of some perfect matchings. In particular, $L(G)=\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$.
@@
A graph $G=(V,E)$ is Birkhoff–von Neumann (BvN) if $P(G)=\left\{x\in\mathbb{R}_{\geq 0}^{E}:x(\delta(v))=1,\,\forall v\in V\right\}$ [4] (see also [2, 8]). That is, $G$ is not BvN if, and only if, $P(G)$ has a facet-defining inequality not exposed by a non-negativity inequality. By the Edmonds-Johnson theorem, such a facet must be exposed by (a special type of) an odd cut, which we refer to as a separating facet-defining cut of $G$. It is well-known that every bipartite graph is BvN, though some non-bipartite graphs such as $K_{4}$ can also be BvN. We shall prove the following lemma.
@@
Let $G=(V,E)$ be a brick, and $d$ the (affine) dimension of $P(G)$. Suppose every face of dimension d2d-2 is exposed by a non-negativity inequality. Then $|V|\leq 10$. Furthermore, either $G$ is the Petersen graph, or $G$ is BvN, or there exist a perfect matching $M$ and a separating facet-defining cut $C$ such that $|M\cap C|=3$.
@@
This is a special case of a more general result by Campos and Lucchesi [3], claiming in particular that in $G$ as above, for every non-tight separating cut $C$, there is a perfect matching $M$ such that $|M\cap C|=3$. The Intersection Theorem is a slight variant of the main theorem in the two papers [5, 6], which establishes in an arbitrary matching-covered graph the minimum intersection size between a perfect matching and a separating cut with intersection size at least three.
@@
In §2, we give a brief overview of necessary notions and results from matching theory, and in §3, we give a short proof of the Petersen Graph Lemma. In §4, we present further preliminary results, and in §5, we prove the Intersection Theorem. Finally, in §6, we prove the Integral Basis Theorem, and then obtain the Main Theorem as a consequence.
2 Ingredients for the Petersen Graph Lemma
Fix a matching-covered graph $G=(V,E)$. For an odd cut $C$ and an edge $e\in E$, denote by $P(G;C)$ and $P_{e}(G)$ the faces $P(G)\cap\{x:x(C)=1\}$ and $P(G)\cap\{x:x_{e}=0\}$ of $P(G)$, respectively.
2.1.
Separating cuts. Let $C=\delta(X)$ be an odd cut such that 1<|X|<|V|11<|X|<|V|-1. We say that $C$ is separating, or contractible, in $G$ if $P(G;C)$ is not contained in $\{x:x_{e}=0\}$ for any edge $e\in E$. We refer to $G/X,G/\bar{X}$ as the $C$-contractions or cut-contractions of $G$. We denote the contraction vertices in a cut-contraction $G/X$ by the corresponding lower case letter $x$. Observe that $C$ is separating if, and only if, both $C$-contractions are matching-covered. Observe further that if $\delta(X),\delta(Y)$ are odd cuts of $G$ such that $X\subset Y$, where $\delta(Y)$ is separating in $G$, and $\delta(X)$ is separating in $G/\bar{Y}$, then $\delta(X)$ is separating in $G$.
$2.$
Bricks. A tight cut is an odd cut $C=\delta(X)$ such that 1<|X|<|V|11<|X|<|V|-1 and $P(G;C)=P(G)$. Clearly, a tight cut is separating. A matching-covered graph with no tight cut is a brick if it is non-bipartite, and is a brace otherwise. Cut-contractions along tight cuts repeatedly give rise to a binary tree rooted at $G$ whose leaves correspond to a tight cut decomposition of $G$ into bricks and braces. The matching polytope $P(G)$ is known to have dimension |E||V|+1b(G)|E|-|V|+1-b(G), where $b(G)$ is the number of bricks of $G$ in a (or any) tight cut decomposition [16, 10]. In fact, the list of bricks and braces in a tight cut decomposition is unique up to the multiplicity of the edges in each brick and brace [13]. Finally, $b(G)=0$ if and only if $G$ is bipartite [16].
2.3.
Separating facet-defining cuts. Let $\delta(X)$ be an odd cut that is facet-defining for $P(G)$. In particular, 1<|X|<|V|11<|X|<|V|-1. Observe that $C$ is not separating if, and only if, $x_{e}\geq 0$ defines the same facet of $P(G)$. Subsequently, $G$ is not BvN if, and only if, there exists a separating facet-defining cut.
@@ 4 Ingredients for the Intersection Theorem
4.5.
Number of bricks in cut-contractions of near-bricks. (1) In general, if one of $G_{1},G_{2}$ is bipartite, then a simple counting argument implies that $C$ is a tight cut. That is, if $C$ is not tight, then $\min\{b(G_{1}),b(G_{2})\}\geq 1$. (2) The converse also holds if $G$ is a near-brick, by 4.4. (3) Subsequently, if $G$ is a near-brick, then $C$ is facet-defining if, and only if, both $G_{1},G_{2}$ are near-bricks.
4.6.
Tight cuts in near-bricks and barriers. Suppose $G$ is a near-brick, and $C$ is a tight cut of $G$. Then one of $G_{1},G_{2}$ is a near-brick while the other one, say $G_{1}$, is bipartite, by 4.4. Let $B$ be the part in a bipartition of $G_{1}$ where $x\notin B$. Observe that $B$ is an independent set of $G$, and $G\setminus B$ has exactly $|B|$ connected components, one of which is $X$, and all others are singletons. We say that $B$ is a barrier in $G$, and $C$ is a barrier cut with barrier $B$. See Figure 2 (left) for an illustrated example. A barrier is maximal if it is not contained in another barrier. [^1] Barriers are defined more broadly and for all matching-covered graphs, but for our purposes, we shall focus only on the (special type of) barriers defined here for near-bricks.
@@
To this end, as $D$ is separating in $G_{2}$, and $C$ is separating in $G$, it follows that $D$ is separating in $G$. As $G_{2}$ is a near-brick, both $G_{2}/Y,G_{2}/\bar{Y}$ are near-bricks, by 4.5, part (3). Thus, $G/\bar{Y}=G_{2}/\bar{Y}$ is a near-brick. As $G_{2}/Y$ is a near-brick, $C=\delta(\bar{X})$ is a tight cut in $G/Y$, and $G_{1}$ is bipartite, it follows that $G/Y$ is a near-brick. Subsequently, both $G/Y,G/\bar{Y}$ are near-bricks, implying that $D$ is a separating facet-defining cut of $G$, by 4.5, part (3).
4.8.
@@
Otherwise, $X_{1},X_{2}$ cross. Let $I:=\delta(X_{1}\cap X_{2})$ and $U:=\delta(X_{1}\cup X_{2})$. Given that $P(G;C_{1})\cap P(G;C_{2})=P(G;C_{1})\not\subseteq\{x:x_{e}=0\}$ for any edge $e\in E$, it follows from 4.2 that $P(G;C_{1})=P(G;I)\cap P(G;U)$. As $P(G;C_{1})$ is a facet of $P(G)$, at least one of $I,U$ must define the same facet as $C_{1}$, say it is $I$. By applying the argument above to $X_{1},X_{1}\cap X_{2}$, and also to $X_{2},X_{1}\cap X_{2}$, we obtain that $C_{1},I$ and $C_{2},I$ are equivalent, so $C_{1},C_{2}$ are equivalent, as claimed.
4.9.
@@
The existence of $M_{0},M_{1},\ldots,M_{d}$ now follows by recursively applying 4.1, part (4).
4.10.

### v3

@@ Abstract
Let $G=(V,E)$ be a matching-covered graph, denote by $P$ its perfect matching polytope, and by $L$ the integer lattice generated by the integral points in $P$. In this paper, we give short, polyhedral proofs for two difficult results established by Lovász (1987), and by Carvalho, Lucchesi, and Murty (2002) in a series of three papers totaling over 120 pages. More specifically, we prove that $L$ has a lattice basis consisting solely of incidence vectors of some perfect matchings of $G$, $2x\in L$ for all $x\in\operatorname{lin}(P)\cap\mathbb{Z}^{E}$, and if $G$ has no Petersen brick then $L=\operatorname{lin}(P)\cap\mathbb{Z}^{E}$. Our proof avoids major technical aspects of the previous proofs, the most important of these being a characterization of the dual lattice, and a ‘Petersen-brick-sensitive’ ear decomposition result for matching-covered graphs. This is achieved by a novel study of the facial structure of the polytope $P$ and its relationship with the lattice $L$. Along the way, we give a new polyhedral characterization of the Petersen graph.
@@ 1 Introduction
Perfect matchings lie at the heart of combinatorial optimization and structural graph theory. Consider a matching-covered graph $G=(V,E)$, that is, a graph where every edge appears in a perfect matching. Denote by $P(G)\subseteq\mathbb{R}^{E}$ the polytope whose vertices are the incidence vectors of the perfect matchings of $G$, and by $L(G)\subseteq\mathbb{Z}^{E}$ the (integer) lattice generated by the integral points in $P(G)$, i.e., $L(G)$ is the set of all integer linear combinations of $P(G)\cap\{0,1\}^{E}$. Understanding the relationship between $P(G)$, the combinatorial structure of $G$, and the algebra of $L(G)$ lies at the center of matching theory.
@@
A seminal result in combinatorial optimization due to Edmonds [10] is that $P(G)$ can be described by non-negativity inequalities: $x_{e}\geq 0,e\in E$, degree equations: $x(\delta(v))=1,v\in V$, and odd cut inequalities amounting to $x(\delta(U))\geq 1$ for all odd-sized subsets $U\subset V$. Here, $x(A)$ denotes $\sum_{a\in A}x_{a}$ for an edge set $A\subseteq E$. Later Seymour [18] gave a graph-theoretic proof of this result using Tutte’s characterization of perfect matchings [19]. In the same paper, Seymour connected the polyhedral study of $P(G)$ with $L(G)$. Through a long and technical argument, he proved that in a bridgeless cubic graph, the all-$2$s vector belongs to $L(G)$, and in fact the all-$1$s vector belongs to the lattice if $G$ has no Petersen minor. This, in turn, established weaker variants of the Berge-Fulkerson conjecture and Tutte’s $4$-flow conjecture, respectively [12, 20].
@@
Many subsequent works have led to a deeper understanding of the structure of $P(G)$ and $L(G)$, as well as various connections between these objects. Naddef [17] and Edmonds, Pulleyblank, and Lovász [11] determined the dimension of $P(G)$ based on structural properties of $G$, showing how combinatorial decomposition manifests itself in the geometry of the polytope. These and related lines of questions led to the development of a fascinating area known as matching theory [15, 13]. In a seminal paper of the area, Lovász [14] gave a far-reaching extension of Seymour’s result and characterized the lattice $L(G)$ itself. He proved that $2x\in L(G)$ for all $x\in\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$, and if $G$ has no Petersen brick (defined later) then $L(G)=\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$. Here, $\operatorname{lin}(S)$ denotes the linear hull of $S$ for $S\subseteq\mathbb{R}^{E}$. The core of his argument was proving a difficult lemma that characterized the dual lattice for a non-Petersen brick. This lengthy proof relied on key tools from matching theory, including the notions of barrier cuts, 2-separation cuts, and ear decompositions.
@@
More recently, Carvalho, Lucchesi, and Murty proved that $L(G)$ has a lattice basis $B$ consisting solely of incidence vectors of some perfect matchings of $G$. That is, $B$ is a (linear) basis for $\operatorname{lin}(L(G))$, and every vector in $\operatorname{lin}(L(G))\cap\mathbb{Z}^{E}$ is an integer linear combination of $B$. This fact was featured as a key application in a series of three excellent papers [6, 7, 8], totaling over 120 pages. There, the authors develop a special type of ear decomposition of the graph $G$ that, vaguely speaking, is sensitive to the number of bricks and Petersen bricks. This in turn gave an affirmative answer to a question previously raised by Murty ([16], Problem 7.3) about the existence of such a basis.
1.1 Main theorem
In this paper, we give relatively short, polyhedral proofs of the theorems mentioned above due to Lovász, and Carvalho, Lucchesi, and Murty. To formally state these theorems, it is necessary to introduce some notation and concepts. Let $G=(V,E)$ be a matching-covered graph. For an odd cut $C$ and an edge $e\in E$, denote by $P(G;C)$ and $P_{e}(G)$ the faces $P(G)\cap\{x:x(C)=1\}$ and $P(G)\cap\{x:x_{e}=0\}$ of $P(G)$, respectively. Note that $P_{e}(G)\subsetneq P(G)$ as $G$ is matching-covered.
Let $C=\delta(X)$ be an odd cut. We say that $C$ is separating, or contractible, in $G$ if 1<|X|<|V|11<|X|<|V|-1, and each $e\in E$ belongs to some perfect matching $M$ of $G$ with $|M\cap C|=1$. The second condition is equivalent to the polyhedral condition that $P(G;C)\not\subseteq P_{e}(G)$ for any edge $e\in E$. Given a separating cut $C=\delta(X)$, we refer to $G/X,G/\bar{X}$ as the $C$-contractions or cut-contractions of $G$. We denote the contraction vertices in a cut-contraction $G/X$ by the corresponding lower case letter $x$. An odd cut $C$ is separating if, and only if, both $C$-contractions are matching-covered. Observe further that if $\delta(X),\delta(Y)$ are odd cuts of $G$ such that $X\subset Y$, where $\delta(Y)$ is separating in $G$, and $\delta(X)$ is separating in $G/\bar{Y}$, then $\delta(X)$ is separating in $G$. This is because any perfect matching of a separating cut-contraction can be extended to a perfect matching of the original graph.
A tight cut is an odd cut $C=\delta(X)$ such that 1<|X|<|V|11<|X|<|V|-1, and every perfect matching of $G$ intersects $C$ exactly once. The second condition is equivalent to the polyhedral condition $P(G;C)=P(G)$. Tight cuts are clearly separating. A matching-covered graph with no tight cut is a brick if it is non-bipartite, and is a brace otherwise. Cut-contractions along tight cuts repeatedly give rise to a binary tree rooted at $G$ whose leaves correspond to a tight cut decomposition of $G$ into bricks and braces. In fact, the list of bricks and braces in a tight cut decomposition is unique up to the multiplicity of the edges in each brick and brace [14].
One specific brick plays an important role in matching theory, namely, the Petersen brick. It can be checked that the Petersen graph (see Figure 1 for an illustration) is a brick. Throughout the work, we allow graphs to have parallel edges, but not loops; with this in mind, we refer to any graph whose underlying simple graph is the Petersen graph, as the Petersen brick.
The main result of this work is a polyhedral proof of the following theorem.
@@ 1.1 Main theorem
Let $G=(V,E)$ be a matching-covered graph, let $L:=L(G)$ and $\bar{L}:=\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$. Then $L$ has a lattice basis consisting solely of perfect matchings of $G$. Furthermore, if $G$ has $p$ Petersen bricks in its tight cut decomposition, then
@@
Similar to the approach of [4] and [8], we make extensive use of tight and more generally separating cuts. However, apart from these, our proof mostly circumvents the lengthy graph-theoretic arguments, and in particular has no dependence on matching theoretic notions such as the uniqueness of the bricks and braces, $2$-separation cuts, braces, removable edges, and ear decompositions; our proof has minimal dependence on the notions of barriers and bricks. It also eliminates the need to study the dual lattice through matching-integral vectors. Instead, our proof is based on polyhedral, or otherwise polyhedrally-driven graph-theoretic arguments that study the relationship between the lattice and the facial structure of the polytope $P(G)$.
@@
Let $G=(V,E)$ be a Petersen-free matching-covered graph. Then $\operatorname{lin}(P(G))$ has an integral basis consisting solely of incidence vectors of perfect matchings. In particular, $L(G)=\operatorname{lin}(P(G))\cap\mathbb{Z}^{E}$.
1.2 Proof ingredients.
@@
A primer on polyhedral theory.
To state the two key ingredients, we first recall necessary definitions from polyhedral theory. We also refer the reader to [3] for more details. Given a polyhedron $Q=\{x\in\mathbb{R}^{n}:Ax\leq b,Cx=d\}$, a face of $Q$ is any nonempty polyhedron obtained by intersecting $Q$ with $\{x:a^{\top}x=\beta\}$ for some inequality $a^{\top}x\leq\beta$ valid for $Q$. Such an inequality $a^{\top}x\leq\beta$ defines, or exposes the face $Q\cap\{x\in\mathbb{R}^{n}:a^{\top}x=\beta\}$. We further classify faces based on a measure called dimension: the (affine) dimension of a polyhedron $Q$ is the maximum number of affinely independent points in $Q$ minus one, denoted by $\dim(Q)$. A facet of $Q$ is a face whose dimension is exactly $\dim(Q)-1$; furthermore, an inequality $a^{\top}x\leq\beta$ of $Ax\leq b$ is facet-defining if it exposes a facet of $Q$. Facets play an essential role in the description of the polyhedron, as any polyhedron is described by its implicit equality constraints (which include $Cx=d$ and possibly a subset of $Ax\leq b$) and its facet-defining inequality constraints (which form a subset of $Ax\leq b$).
@@
Fix a matching-covered graph $G=(V,E)$. The matching polytope $P(G)$ is known to have dimension |E||V|+1b(G)|E|-|V|+1-b(G), where $b(G)$ is the number of bricks of $G$ in a tight cut decomposition [17, 11]. Further, $b(G)=0$ if and only if $G$ is bipartite [17].
As stated above, the perfect matching polytope $P(G)$ can be described by three types of constraints: degree constraints, non-negativity constraints, and odd cut constraints. Since the degree constraints are equalities, there are only two types of facets in the description of $P(G)$, exposed by (i) $x_{e}\geq 0$ for some $e\in E$, or (ii) $x(C)\geq 1$ for some odd cut $C\subseteq E$. Note that in general, the facets of $P(G)$ might be exposed by both types of inequalities at once. In this work, we distinguish a class of graphs where each facet is exposed by some non-negativity constraint. Formally, a graph $G=(V,E)$ is Birkhoff–von Neumann (BvN) if $P(G)=\left\{x\in\mathbb{R}_{\geq 0}^{E}:x(\delta(v))=1,\,\forall v\in V\right\}$. This terminology was coined recently in [5], but the class was studied and characterized much earlier, see [2, 9]. It is well-known that every bipartite graph is BvN, though some non-bipartite graphs such as $K_{4}$ can also be BvN.
Suppose a matching-covered graph $G$ is not BvN. By definition, this happens if, and only if, $P(G)$ has a facet that is not exposed by any non-negativity inequality $x_{e}\geq 0$. Such a facet must be exposed by (a special type of) an odd cut inequality, denote it $x(C)\geq 1$. Observe that $C$ must be separating since otherwise, there would exist an edge $e\in E$ such that $P(G)\cap\{x:x(C)=1\}\subseteq P(G)\cap\{x:x_{e}=0\}\subsetneq P(G)$ — recall that $G$ is matching-covered. However, as $x(C)\geq 1$ exposes a facet of $P(G)$, the inequality $x_{e}\geq 0$ must expose the same facet, contradicting our choice of the facet. Hence, such a cut $C$ must be separating; we refer to $C$ as a separating facet-defining cut of $G$. Consequently, $G$ is not BvN if, and only if, there exists a separating facet-defining cut.
We signify the role of separating facet-defining cuts in the following theorem, which also provides a novel polyhedral characterization of the Petersen graph, distinguishing it from all other bricks.
@@
Let $G=(V,E)$ be a brick, and $d$ the dimension of $P(G)$. Suppose every face of dimension d2d-2 is exposed by a non-negativity inequality. Then $|V|\leq 10$. Furthermore, $G$ is the Petersen graph, or $G$ is BvN, or there exist a perfect matching $M$ and a separating facet-defining cut $C$ such that $|M\cap C|=3$.
@@
This is a special case of a more general unpublished result by Campos and Lucchesi [4], claiming in particular that in $G$ as above, for every non-tight separating cut $C$, there is a perfect matching $M$ such that $|M\cap C|=3$. The Intersection Theorem is a slight variant of the main theorem in the two papers [6, 7], which establishes in an arbitrary matching-covered graph the minimum intersection size between a perfect matching and a separating cut with intersection size at least three.
@@
In §2, we give a short proof of the Petersen Graph Lemma. In §3, we present further preliminary results, and in §4, we prove the Intersection Theorem. Finally, in §5, we prove the Integral Basis Theorem, and then obtain the Main Theorem as a consequence.
@@ 3 Ingredients for the Intersection Theorem
3.5.
Number of bricks in cut-contractions of near-bricks. (1) In general, if one of $G_{1},G_{2}$ is bipartite, then a simple counting argument implies that $C$ is a tight cut. That is, if $C$ is not tight, then $\min\{b(G_{1}),b(G_{2})\}\geq 1$. (2) The converse also holds if $G$ is a near-brick, by 3.4. (3) Subsequently, if $G$ is a near-brick, then $C$ is facet-defining if, and only if, both $G_{1},G_{2}$ are near-bricks.
3.6.
Tight cuts in near-bricks and barriers. Suppose $G$ is a near-brick, and $C$ is a tight cut of $G$. Then one of $G_{1},G_{2}$ is a near-brick while the other one, say $G_{1}$, is bipartite, by 3.4. Let $B\subseteq V$ be the part in a bipartition of $G_{1}$ where the shrunk vertex $x\notin B$. Observe that $B$ is an independent set of $G$, and $G\setminus B$ has exactly $|B|$ connected components, one of which is $X$, and all others are singletons. We say that $B$ is a barrier in $G$, and $C$ is a barrier cut with barrier $B$. See Figure 2 (left) for an illustrated example. A barrier is maximal if it is not contained in another barrier. [^1] Barriers are defined more broadly and for all matching-covered graphs, but for our purposes, we shall focus only on the (special type of) barriers defined here for near-bricks.
@@
To this end, as $D$ is separating in $G_{2}$, and $C$ is separating in $G$, it follows that $D$ is separating in $G$. As $G_{2}$ is a near-brick, both $G_{2}/Y,G_{2}/\bar{Y}$ are near-bricks, by 3.5, part (3). Thus, $G/\bar{Y}=G_{2}/\bar{Y}$ is a near-brick. As $G_{2}/Y$ is a near-brick, $C=\delta(\bar{X})$ is a tight cut in $G/Y$, and $G_{1}$ is bipartite, it follows that $G/Y$ is a near-brick. Subsequently, both $G/Y,G/\bar{Y}$ are near-bricks, implying that $D$ is a separating facet-defining cut of $G$, by 3.5, part (3).
3.8.
@@
Otherwise, $X_{1},X_{2}$ cross. Let $I:=\delta(X_{1}\cap X_{2})$ and $U:=\delta(X_{1}\cup X_{2})$. Given that $P(G;C_{1})\cap P(G;C_{2})=P(G;C_{1})\not\subseteq\{x:x_{e}=0\}$ for any edge $e\in E$, it follows from 3.2 that $P(G;C_{1})=P(G;I)\cap P(G;U)$. As $P(G;C_{1})$ is a facet of $P(G)$, at least one of $I,U$ must define the same facet as $C_{1}$, say it is $I$. By applying the argument above to $X_{1},X_{1}\cap X_{2}$, and also to $X_{2},X_{1}\cap X_{2}$, we obtain that $C_{1},I$ and $C_{2},I$ are equivalent, so $C_{1},C_{2}$ are equivalent, as claimed.
3.9.
@@
The existence of $M_{0},M_{1},\ldots,M_{d}$ now follows by recursively applying 3.1, part (4).
3.10.